OpenStudy (gabylovesyou):

Given a polynomial f(x), if (x − 1) is a factor, what else must be true? f(0) = 1 f(1) = 0 f(−1) = 0 f(0) = −1

OpenStudy (gabylovesyou):

@mathmale @hartnn @ParthKohli

hartnn (hartnn):

x-1 is a factor, which means if we put x= 1, we get f(x) = 0

hartnn (hartnn):

which is same as saying f(1) = 0

OpenStudy (damonte212):

If x=1, you get f(1)=(1-1)*g(1)=0*g(1)=0 So, you get that f(1)=0 is true.

OpenStudy (gabylovesyou):

so B... Thank you so much :)